# Discrete Heat Kernels on Simplicial Complexes and Its Application to Functional Brain Networks

### v1

@@ 4 Discussion and conclusion
Simplicial diffusion offers a simple, topology–aware regularizer for functional networks. Applied to the MTA group averages, it suppresses isolated noisy connections and sharpens coherent patterns without imposing modular structure or hard thresholds. Relative to unsmoothed means, the smoothed contrasts between users and non–users are more spatially contiguous and anatomically interpretable. Because the heat kernel operator preserves the simplex-space topology, the procedure does not create spurious $k$-cycles and can be coupled with downstream topological summaries (e.g., loop/non–loop decompositions) for further analysis [2]. Limitations include the choice of bandwidth $t$ and the reliance on a single atlas. In practice $t$ can be selected by minimizing reconstruction error on held–out simplices or by inspecting stability across a small grid; multi–atlas analyses can assess robustness. The present demonstration uses group means, future work would explore subject-level smoothing with appropriate statistical frameworks for group comparisons, potentially using permutation testing on smoothed edge weights or topological features. Integration with dynamic connectivity analyses could reveal how cannabis affects the temporal evolution of network states, with heat kernel smoothing potentially stabilizing time-varying connectivity estimates [13, 16]. Overall, simplex smoothing improves signal-to-noise in functional connectomes with minimal assumptions, producing clearer summaries for visualization and group comparison.

### v2

@@ 4 Discussion and conclusion
Simplicial diffusion offers a simple, topology–aware regularizer for functional networks. Applied to the MTA group averages, it suppresses isolated noisy connections and sharpens coherent patterns without imposing modular structure or hard thresholds. Relative to unsmoothed means, the smoothed contrasts between users and non–users are more spatially contiguous and anatomically interpretable. Because the heat kernel operator preserves the simplex-space topology, the procedure does not create spurious $k$-cycles and can be coupled with downstream topological summaries (e.g., loop/non–loop decompositions) for further analysis [anand2021hodge]. Limitations include the choice of bandwidth $t$ and the reliance on a single atlas. In practice $t$ can be selected by minimizing reconstruction error on held–out simplices or by inspecting stability across a small grid; multi–atlas analyses can assess robustness. The present demonstration uses group means, future work would explore subject-level smoothing with appropriate statistical frameworks for group comparisons, potentially using permutation testing on smoothed edge weights or topological features. Integration with dynamic connectivity analyses could reveal how cannabis affects the temporal evolution of network states, with heat kernel smoothing potentially stabilizing time-varying connectivity estimates [dakurah2025robust, dakurah2024subsequence]. Overall, simplex smoothing improves signal-to-noise in functional connectomes with minimal assumptions, producing clearer summaries for visualization and group comparison.